OpenStudy (anonymous):

Solve the absolute value equation. Graph the solution. 2|3x-5| - 8 = 8 A. X=-1 5/6 or x = 5/6 B. x = -1 or x = 4 1/3 C. x = -1 5/6 or x = 4 1/3 D. x =-1 5/6 or x = -1

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Preetha

OpenStudy (wolf1728):

2|3x-5| - 8 = 8 dividing by 2 |3x -5| -4 = 4 |3x -5| = 8 X = 4 1/3 Check 2|3x-5| - 8 = 8 2 * |3x -5| -8 = 8 2 * |13 -5| -8 = 8 16 = 16

OpenStudy (anonymous):

so its b?

OpenStudy (wolf1728):

Yes

OpenStudy (anonymous):

thank you it was helpful

OpenStudy (wolf1728):

Glad to help